If we have given three resistances, current and voltage...how we find the 4th resistance


1
Expert's answer
2021-03-10T17:16:49-0500

Answer

According to physics

Wheatstone bridge is an electrical circuit used to measure an unknown  resistance by balancing two legs of a bridge circuit, one leg of which includes the unknown resistance.




Wheatstone bridge circuit diagram. The unknown resistance Rx is to be measured; resistances R1R2 and R3 are known, where R2 is adjustable. When the measured voltage VG is 0, both legs have equal voltage ratios: "\frac{R_2}{R_1} = \frac{R_x}{R_3} and\\ R_x=\frac{ R_3}{R_2}{R_1}"